Suppose marginal utility of good X is 20 while its price is Rs. 4 per unit and marginal utility of Y good is 50 while its price is Rs.5 per unit .The individual to whom this information applies is spending 20 on each good .Is he maximizing his satisfaction.?





(12+6.75)





(i) eku yhft, fd oLrq X dh lhekar mi;ksfxrk 20 gS tcfd bldh dher

1
Expert's answer
2021-12-03T11:34:26-0500

Task #274645


Solution


To maximize total utility for every dollar that the individual spends , they should spend it on the item which yields the greatest marginal utility per dollar of expenditure.


The individual will only get maximum satisfaction if they attain the highest level of satisfaction from their economic decisions.


The maximum satisfaction will be attained at a point where the marginal utility per dollar for good X is equal to the marginal utility per dollar for good Y.


Marginal utility per dollar is the amount of additional utility the individual receives given the price of the product.


Where the marginal utility per dollar is calculated as follows:


"\frac{MU}{P }"


To maximize utility :


"\frac{Px}{Py}=\frac{MUx}{MUy}"

Or


"\frac{MUx}{Px}=\frac{MUy}{Py}"


Marginal utility per dollar for X:

"\frac{20}{4}=5"


Marginal rate per dollar for Y:


"\frac{50}{5}=10"


Comparing the two goods, marginal utility per dollar for good X and Y are not the same such that the marginal utility per dollar for good Y is twice that of X.


If the individual spends 20 on each good he or she will have 5 units of good X and 4 units of good Y which will not be enough to attain maximum satisfaction.


The individual will thus not attain maximum satisfaction unless the combination of goods gives the same marginal utility per dollar for X and Y at the chosen point on the budget line.